OpenStudy (anonymous):

Find X^2 + y^2=1 intersection with y=x

OpenStudy (amistre64):

+- (cos(45),sin(45))

OpenStudy (anonymous):

ohh so did good

OpenStudy (amistre64):



OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yea but i have to show how i got it.. so i subtitute x^2+(x)^2=1 and solve and i got the answer for x then i did the same thing for y

hero (hero):

Pssh....that was easy

myininaya (myininaya):

you could do it that way but amistre was showing you how to get the answer graphically/geometrically

OpenStudy (anonymous):

yea,,, i agree.. but i want to check if what i did after the graph was ok

myininaya (myininaya):

\[x^2+x^2=1 =>2x^2=1=> x^2=\frac{1}{2} => x=\pm \sqrt{\frac{1}{2}} =\pm \frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

yeaaaa:)

hero (hero):

I was just about to do that myininaya is too quick

myininaya (myininaya):

\[x=y => y=x=\pm \frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

i have to do it both way,,, so yea that was my answer...

myininaya (myininaya):

\[(\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2}) or (-\frac{\sqrt{2}}{2}, -\frac{\sqrt{2}}{2})\]